A restaurant offers a 1kg steak as part of a promotion. Assume weights of the steaks are approximately normally distributed with a standard deviation of 25 grams.
X= weights of the steaks in grams
X bar= average weights of the steaks in grams
Assume that the average weights on 5 restaurants has been investigated.

Calculate the 75th percentile of the average weight of steaks for the 5 restaurants.
1
Expert's answer
2020-10-01T15:15:58-0400

solution


let "X_{0.75}" be the weight of the steak at the 75th percentile. "75\%" of the steaks will weigh less than "X_{0.75}" . Mean weight of steak served "\bar X =1\ kg =1000\ grams" and standard deviation "\sigma =25"


The corresponding "Z- value" is

"0.67"



"Z=\frac{X_{0.75}-\bar X}{\sigma}"

"0.67=\frac{X_{0.75}-1000}{25}"

"X_{0.75}=0.67*25+1000 = 1016.75"

answer: the 75th percentile of the average weights is 1016.75 grams or 1.01675 kg